OpenStudy (landon34):

One number has a prime factorization of 2 3 • 3 2, and another number has a prime factorization of 2^2 • 3^3. Which of the following expressions would equal the greatest common factor of these two numbers? 2^5 • 3^5 2^3 • 3^3 2^2 • 3^2 2 • 3

OpenStudy (anonymous):

one is \(2^2\times 3^2\) and the other is \(2^2\times 3^3\) like that?

OpenStudy (landon34):

yes.

OpenStudy (anonymous):

take each factor you see to the lowest power you see in any one number in this case the lowest power of 3 is 2, so it would be \[2^2\times 3^2\]

OpenStudy (landon34):

2^3

OpenStudy (landon34):

the first number is 2^3 instead of 2^2